OpenStudy (anonymous):

help me. 6y+2y=16. you have to solve it

jigglypuff314 (jigglypuff314):

you can add together numbers that have the same variable attached to them so like 3y + 5y = 8y so do you know what 6y + 2y = ? :)

OpenStudy (anonymous):

yeah its 8y but how am i supposed to solve it? @jigglypuff314

jigglypuff314 (jigglypuff314):

so now you have 8y = 16 right? 8 is attached to the y by multiplication to separate them you would have to divide both sides by 8

OpenStudy (anonymous):

i dont know what my teacher means by solve it it is already solved. 6y+2y=16. but 6y+2y=8y not just plaine 16. im sooo confused!

OpenStudy (anonymous):

oooooooooooooooooooooooo i get it now!!!!!!!! hahahahahahaha thank you!!!!!!!

jigglypuff314 (jigglypuff314):

you're welcome :)